OpenStudy (anonymous):

Which of the following sets, if any, is closed under multiplication? a) {0,3} b) {1} c) {6,12,18,24...}

OpenStudy (zzr0ck3r):

only {1} a set is closed iff any product of elements is in the set.. so 1*1 = 1*1 = {1} but 3*3 = 9 is not in the set {0,3} and 12*12 = 144 is not in the last set

OpenStudy (anonymous):